The half-life of 23191Pa is 3.25 × 104 y. How much of an initial 10.40 microgram sample remains after 3.25 × 105 y?

Answer is 0.0102
Answer #1

Given that the half life of 231Pa91 is 3.25 x 104 year

Initial amount is 10.4 x 10-6 grams, t = 3.25 x 105 year

we can use the integrated rate law for a first-order nuclear reaction

That is ln(N/No) = -kt and k = 0.693/t1/2

ln(N/10.4*10-6) = - (0.693/3.25*104)*(3.25*105) -------------------------------(1)

The above equation (1) gives N = 1.02 x 10-8 grams of 231Pa91

Therefore, the sample remaining after 3.25 x 105 years is 0.0102 Micrograms